OpenStudy (anonymous):

Find the standard form of the equation of a parabola of the form y=a(x-h)^2+k whose vertex is (2,1) and passes through the point (3,2)

OpenStudy (yttrium):

What do you understand about standard form @Tiara89 ?

OpenStudy (anonymous):

absolutely nothin its an online math class and i regret it! @Yttrium

OpenStudy (yttrium):

Well. It's just a form of equation that is already expanded and equated to 0.

OpenStudy (yttrium):

Let's solve first for a. Since we have given the vertex in form (h,k) --> (h=2,k=1) and one point at (x=3,y=2) Hence we can solve for a.

OpenStudy (yttrium):

@Tiara89 , can you find a?

hero (hero):

Input the vertex and the point in to the equation and solve for a: 2=a(3-2)^2+1 2 = a + 1 a = 1 Now you can write the equation in vertex form: y = (x - 2)^2 + 1 To write in standard form, simply expand the right side: y = (x - 2)(x - 2) + 1 y = x(x - 2) - 2(x - 2) + 1 y = x^2 - 2x - 2x + 4 + 1 y = x^2 - 4x + 5 Remember, the standard form of a quadratic equation is y = ax^2 + bx + c